Coffee worth $ 1.05 per pound is mixed with coffee worth $0.85 per pound to obtain 20 pounds of mixture worth $0.90 per pound. How many pounds of each type of coffee were used?
Answer by stanbon(75887)

Coffee worth $ 1.05 per pound is mixed with coffee worth $0.85 per pound to obtain 20 pounds of mixture worth $0.90 per pound. How many pounds of each type of coffee were used?
----------------------
Equation:
value + value = value
1.05x + 0.85(20-x) = 0.90*20
---------------------------------
105x + 85*20 - 85x = 90*20
20x = 5(20)
x = 5 lbs (amt. of $1.05 coffee needed)
-----
20-x = 15 lbs (amt. of $0.85 coffee needed)
---------------
Cheers,
Stan H.
---------------